OpenStudy (mathhelp346):

How can I determine whether or not m,n,sqrt{m^2+n^2} is a Pythagorean triple or not?

OpenStudy (anonymous):

if this triple satisfies the equation \(a^2+b^2=c^2\) then its a pyth triple

OpenStudy (anonymous):

go ahead :)

OpenStudy (mathhelp346):

but how do u find the square root of m^2+n^2?

OpenStudy (anonymous):

we dont need to find square root of m^2+n^2 because we dont know what numbers are them

OpenStudy (anonymous):

its a pyth triple because if we let a=m b=n and c=sqrt{m^2+n^2} u have m^2+n^2=(sqrt{m^2+n^2} )^2=m^2+n^2

OpenStudy (mathhelp346):

so then its not a triple because m^2+n^2 dont equal the square root of m^2+n^2

OpenStudy (mathhelp346):

wait what

OpenStudy (anonymous):

3,4,5 is a pyth triple right?

OpenStudy (mathhelp346):

yea

OpenStudy (anonymous):

why?

OpenStudy (mathhelp346):

because 3^2+4^2=5^2

OpenStudy (anonymous):

yup now see if thats right for m,n,sqrt{m^2+n^2} or not?

OpenStudy (mathhelp346):

ohhh i see now

OpenStudy (mathhelp346):

thanks so much

OpenStudy (anonymous):

welcome :)